A source emitting light of wavelengths 480 nm and 600 nm is used in a double slit interference experiment. The separation between the slits is 0.25 mm and the interference is observed on a screen placed at 150 cm from the slits. Find the linear separation between the first maximum (next to the central maximum) corresponding to the two wavelengths.

Grade:11

1 Answers

Kevin Nash
9 years ago
We know that, the first maximum (next to central maximum) occurs at y = λD/d Given that, λ base 1 = 480 nm, λ base 2 = 600 nm, D = 150 cm 1.5 m and d = 0.25 mm = 0.25 * 10^-3 So, y base 1 = Dλ base 1/d = 1.5 * 480 * 10^-9/0.25 * 10^-3 = 2.88 mm Y base 2 = 1.5 * 600 * 10^-9/0.25 * 10^-3 = 3.6 mm. So, the separation between these two bright fringes is given by, ∴ separation = y base 2 – y base 1 = 3.60 – 2.88 = 0.72 mm.